OpenStudy (anonymous):

1. ( 8 + square root 5) (8- square root 5) 2. Create your own binomial expression with a radical in the second term. Part 1: Identify its conjugate and explain, in complete sentences, why it is the conjugate. Part 2: Multiply your original binomial expression and its conjugate. What happened to the radicals and why?

OpenStudy (anonymous):

did you do #1?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

do you know how to FOIL it out?

OpenStudy (anonymous):

I'm having trouble with that :/

OpenStudy (anonymous):

1. Use the property -> (a+b)(a-b) = a^2 - b^2 .. 2. Radical in the second term is -- sqrt 5 .. So and expression using it could be .. \[3 + \sqrt 5\] ... (you may make anyone of your choice..) (i) conjugate of 3+ sqrt 5 = 3 - sqrt 5 (ii) When you multiply these two, you get .. \[(2+\sqrt 5)(2-\sqrt 5)\] \[=> 3^2 - (\sqrt5)^2 = 9-5 = 4\]

OpenStudy (anonymous):

oh, I should let you know that each problem is NOT related to each other in any way, they're different

OpenStudy (anonymous):

would the answer to the first question be 64 - 18 square root 5 ???

OpenStudy (anonymous):

they sort of are related because your second question talks about the conjugate of a binomial... the conjugate of (a+b) is (a-b).... and vice/versa...

OpenStudy (anonymous):

well yes, but I mean they aren't suppose to go together if you know what I mean. And for the first question, is my answer correct??

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

use the formula @Aditi_Singh , stated: \(\large (a+b)(a-b) = a^2 - b^2 \)

OpenStudy (anonymous):

multiplying conjugates will always give you difference of squares....

OpenStudy (anonymous):

\(\large (8+\sqrt5)(8-\sqrt5)=8^2-(\sqrt5)^2 \)

OpenStudy (anonymous):

ohh. when I re-did it, I got 59- 16 square root of 5, idk if thats right...

OpenStudy (anonymous):

\(\large (8+\sqrt5)(8-\sqrt5)=8^2-(\sqrt5)^2=64-5 \)

OpenStudy (anonymous):

=59

OpenStudy (anonymous):

oh it is 59, I did it that way the first time and wasn't sure if it was right bc I didn't get a square root

OpenStudy (anonymous):

now to do the second part do the same thing... change the 8 to any number you want.. and change that \(\sqrt5 \) to any radical you want... then do the same thing we just did on the one you created....:)

OpenStudy (anonymous):

Believe me or not.. but i really think that the 1st and the 2nd ques are related to each other.. in some way!

OpenStudy (anonymous):

thanks guys, I just figured out how to do the second one lol. and they kinda are related loll